OpenStudy (anonymous):

if 289=17^1/5x then x=?

OpenStudy (yttrium):

is the equation like this \[289=17^{1/5x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hey ! What is the question???????????

OpenStudy (anonymous):

289=17power 1/5x

OpenStudy (anonymous):

Oh ! OK !

OpenStudy (yttrium):

\[289 = 17^{1/5x}\] \[289 = 17^{2}\] \[17^{2} = 17^{1/5x}\] What you need to do is to equate the exponents since they have the same base of 17 Can you do it now?

OpenStudy (anonymous):

@:Yttrium : You re falllllllllllllllllllllllllllllllllllllllllllll

OpenStudy (anonymous):

s0 10 is the answer

OpenStudy (anonymous):

NO :)

OpenStudy (yttrium):

1/10, I guess

OpenStudy (anonymous):

NO :)

OpenStudy (yttrium):

10x = 1, Therefore x = 1/10

OpenStudy (yttrium):

@jimra Did you get it?

OpenStudy (anonymous):

i also get the same answer .but is it right?

OpenStudy (yttrium):

Yes, you can check that by substituting the value of x. Check whether the equation will yield to 289 = 289

OpenStudy (anonymous):

right

OpenStudy (anonymous):

HEY look ! First answer this : 17^x=289 What is the x ?!

OpenStudy (yttrium):

2

OpenStudy (zzr0ck3r):

ahh the x is in the denominator...

OpenStudy (yttrium):

Yes.

OpenStudy (anonymous):

Ok ! Now 2/1.5 =? ?=x !!!

OpenStudy (anonymous):

x in numerator

OpenStudy (yttrium):

if x in numerator, therefore 2 = x/5 Hence x = 10 Maybe, i just misinterpret the eq'n. I'm sorry

OpenStudy (zzr0ck3r):

\[17^2=17^{\frac{x}{5}}\]????

OpenStudy (anonymous):

Oh ! OK ! SAY : Your answer is right ! 2.5=10

OpenStudy (zzr0ck3r):

\[17^2=17^{\frac{1}{5}x}\\log_{17}(17^2)=log_{17}{(17^{\frac{1}{5}x}})\\2=\frac{1}{5}x\\x=10\]

OpenStudy (anonymous):

@ Yuttrium : You are true ! @ jimra : The answer is 10 . !!!

OpenStudy (zzr0ck3r):

i thought the answer was 2.5 = 10

OpenStudy (yttrium):

@jimra, Did you get the concept?

OpenStudy (anonymous):

@jimra , Answer to Yttrium question !

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."